OpenStudy (anonymous):

What is the oblique asymptote of the function f(x) = the quantity of x squared plus x minus 2, all over x plus 1? y = x + 3 y = x y = x - 1 y = x - 3 @satellite73

OpenStudy (anonymous):

i guess using the equation editor is out of the question huh? \[\frac{x^2+x-2}{x+1}\]

OpenStudy (anonymous):

should i use it?

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

no no i got it the "all over" and the "plus" is kind of cute

OpenStudy (anonymous):

lol ur soooo funnyy omg u always make my day when im most upset lol

OpenStudy (anonymous):

what you need to do for this one is divide synthetic division is easiest, long division is annoying

OpenStudy (anonymous):

yeah i got it but idk which one is the answer http://www.wolframalpha.com/input/?i=%28x%5E2%2Bx-2%29%2F%28x%2B1%29

OpenStudy (anonymous):

ok let me show you how to interpret the answer you got from wolf

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

you see where it says "quotient and remainder" and has \[x^2+x-2 = x × (x+1)+-2\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

oops it has \[x^2+x-2=x(x+1)-2\] that means \[\frac{x^2+x-2}{x+1}=x-\frac{2}{x+1}\]

OpenStudy (anonymous):

is it y = x-3?

OpenStudy (anonymous):

your slant asymptote is therefore just \(y=x\)

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

no it is just he quotient, ignore the remainer

OpenStudy (anonymous):

oh gotcha okayy!

OpenStudy (anonymous):

we can do another if you like, make sure you know how to get to the answer from the wolf

OpenStudy (anonymous):

yeah that'll be great I always cant seem to get the answer off of wolf cuz it's not that obvious for mw

OpenStudy (anonymous):

you got one or you want me to make one up better if you have one, kill two birds

OpenStudy (anonymous):

LOL okay i'll look for one

OpenStudy (anonymous):

ok i will wait patiently

OpenStudy (anonymous):

okay i have one but instead of asking for the oblique asymptote this one is asking for a vertical asymptotes is that okay?

OpenStudy (anonymous):

yeah but that is way easier no wolf needed because you dont have to divide anything

OpenStudy (anonymous):

go ahead we can do it in five seconds probably with our eyeballs depending on the question

OpenStudy (anonymous):

yeah u have to divide in my question

OpenStudy (anonymous):

ok post !

OpenStudy (anonymous):

What are the vertical asymptotes of the function f(x) = the quantity of 5 x plus 5, all over x squared plus x minus 2?

OpenStudy (anonymous):

you crack me up

OpenStudy (anonymous):

\[f(x)=\frac{5x+5}{x^2+x-2}\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

the quantity of 5 x plus 5, all over x squared plus x minus 2?

OpenStudy (anonymous):

no u said u crack me up lol

OpenStudy (anonymous):

it is cute, very literal i like it in any case lets do this quickly, you want the VERTICAL asymptotes right?

OpenStudy (anonymous):

LOL okay yeah

OpenStudy (anonymous):

we do this by setting the denominator equal to zero and solving for \(x\)

OpenStudy (anonymous):

\[x^2+x-2=0\\ (x+2)(x-1)=0\\ x+2=0,x=-2\\ x-1=0,x=1\]

OpenStudy (anonymous):

two vertical asymptotes are \(=-2\) and \(x=1\)

OpenStudy (anonymous):

i got x = 1 and x = -2

OpenStudy (anonymous):

YAYYYYYY!!!

OpenStudy (anonymous):

i got it *victory dance* oh yeaahhh! LOL

OpenStudy (anonymous):

I was able to find that on wolf!!!!! :D

OpenStudy (anonymous):

cheater

OpenStudy (anonymous):

LOL jellyyy!

OpenStudy (anonymous):

i meant CONGRATULATIONS

OpenStudy (anonymous):

xD THANK YOUUUU!!

OpenStudy (anonymous):

do i get a medal? no? no? ok xD

OpenStudy (anonymous):

yw you done?

OpenStudy (anonymous):

ummmm... idk

OpenStudy (anonymous):

you don't know if you are done??!

OpenStudy (anonymous):

xD IKR.. kinda crazyy but ik im not done it's just that i dont know if i will need u now cuz if i can do others by myself no need to bother u but if im stuck i'll surely bother YA!! hahah

OpenStudy (anonymous):

ok i will do some actual work now if you need help just post

OpenStudy (anonymous):

Okayyy!! gracias!

OpenStudy (anonymous):

ATTENTIONNNNNN!!!!!

OpenStudy (anonymous):

I NEED HELP RIGHT AWAAAYYY!

OpenStudy (anonymous):

zzzzz

OpenStudy (anonymous):

lol WAKE UP!

OpenStudy (anonymous):

i think i have the answer to this one but not sure.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Steve can paint a room three times as fast as Billy. When they work together, Steve and Billy can paint a large room in 4 hours. How many hours would it take Billy to paint it by himself? sixteen-thirds <----- MY ANSWER three sixteenths 16 3

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hold on a sec

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

one rate x, the other 3x

OpenStudy (anonymous):

for a total of 4x

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

4 HOWAAAS

OpenStudy (anonymous):

so \[4x\times 4=1\] making ' \[x=\frac{1}{16}\] and \[3x=\frac{3}{16}\]

OpenStudy (anonymous):

oh c'moooon I FLIPPED THEMM!! lol

OpenStudy (anonymous):

i'm soo dumb LOL

OpenStudy (anonymous):

no it is ok, we still have to flip one is 16 hours the other is 16/3 hours

OpenStudy (anonymous):

but you want the time for the slower one, so i guess you are a bit dim

OpenStudy (anonymous):

try 16

OpenStudy (anonymous):

lol so is it 16/3 or 3/16 cuz i'm confused

OpenStudy (anonymous):

it it is neither because you want the time for the SLOWER one not the faster one faster one's rate is \(\frac{3}{16}\) takes \(\frac{16}{6}\) hours but the SLOWER ones' rate is\[\frac{1}{16}\] takes \[16\]hours

OpenStudy (anonymous):

go with 16

OpenStudy (anonymous):

oh wooow how REATRDED xD

OpenStudy (anonymous):

yeah just what i was thinking reatrded

OpenStudy (anonymous):

OMGG!! LOL noooooooooo xD i swear i can spell LOL

OpenStudy (anonymous):

i can spel to

OpenStudy (anonymous):

now are you done? i got stuff to do

OpenStudy (anonymous):

you should definitely enter a spelling bee LOL

OpenStudy (anonymous):

um no u can go though i'll just get help from someone else.

OpenStudy (anonymous):

ok if i am here i will help

OpenStudy (anonymous):

ok :)